# Optimal Virtual Model Control for Robotics: Design and Tuning of Passivity-Based Controllers

### v2

@@ II-C Virtual structures and implementation
Figure 5: Virtual model controller block diagram. $({\bm{F}}_{e},{\bm{z}}_{e})$ is an external port. $r$ represent external inputs to the controller.
@@
Constructing the Jacobian is a large part of the computational cost for robots/virtual structures with several joints. Explicit computation of the Jacobian can be avoided for serial robots or tree-structured robots by recursive algorithms, which propagate kinematics/velocities forwards through the structure from the base, and forces backward from the tip(s), while also computing the required joint actuation ${\bm{u}}$. For an example, refer to the the backwards pass of the Recursive Newton Euler Algorithm in [47].
@@
Although we limit our discussion to fully actuated robots, virtual model control is still possible for underactuated robots. The mechanism must in that case be designed such that the un-actuated joints are not required to emulate the virtual forces on the robot. Indeed, in Figure 2, the force ${\bm{F}}_{1}$ does not require any actuation of the third joint. Likewise, in a quadruped, the underactuation is due to the lack of rigid connection to the ground. However, one could design a virtual mechanism acting between any part of the body/legs of the robot [19].
III The design of virtual mechanisms
III-A Virtual model control for reaching
The advantages of designing a controller as a virtual mechanisms can be illustrated by revisiting the classical control of a ‘reach’ motion, the basic primitive of any manipulator. Once gravity is compensated, the simplest virtual mechanism is given by a virtual spring and a virtual damper connected between the end effector and the target. This corresponds to the classical PD control in the task space [46]. It is straightforward to see that the adoption of nonlinear elements in this context enables important features. Consider a spring and a damper with nonlinear force-displacement characteristics
@@ III-A Virtual model control for reaching
The more advanced exemplary virtual mechanism of Figure 6 allows for refined control over smoothness and peak velocity of the reaching motion. Linear spring and dampers are now attached between the end effector and virtual moving cart. The cart is pulled by a constant force, experiences some drag due to friction, and has mass, which leads to a dynamic controller.
This design has several useful features. It separates the tasks of constraining the robot and driving motion along the line: the spring-damper components between the damper and the cart regulate the constraint, while forces/components applied to the cart regulate the reaching motion. By driving the cart with a constant force rather than as a function of time, we achieve a safer form of path following. Consider an obstruction such as a person in the path of the robot: the end-effector and person collide and the cart continues to move forwards only until equilibrium is reached between the driving force and the constraint-spring. Thus, even for a stiff spring ensuring a strong constraint to the path, we can achieve compliant behaviour along the direction of motion.
The velocity can be regulated by tuning the friction applied to the cart: neglecting the robot’s friction, the steady state velocity can be determined by balancing the cart’s friction with the applied driving force. Similarly, fast acceleration can be regulated by increasing the inertance of the cart. The cart also has the effect of filtering the driving force: a step change in driving force results in smooth driving torques. For a numerical example of these effects, see Figure 6.
The above considerations illustrate how, in contrast to control synthesis methods based on mathematical derivations, the design of virtual mechanisms carries significant insights on the control action. By relying on physical intuition, we can make predictions about the robot reactions to obstacles, the velocity peak of the end effector, and the filtering effects of the controller dynamics. It is also easier to understand how the parameters of the virtual mechanism will affect the control performance. For example, it is straightforward to imagine that a larger cart mass will induce a stronger lowpass effect on the pulling force, as shown in Figure 6. Similarly, a larger spring stiffness may lead to higher peak forces.
@@
III-B Virtual model control for laparoscopic surgery
Virtual model control appears particularly effective on complex tasks. Here we consider robotic surgery, specifically laparoscopic surgery, as shown in Figure 7. The controller has two main objectives: (i) the instrument must pass through a small aperture in the patient’s body, often achieved by maintaining the tool over the so-called Remote Center of Motion (RCM), a fixed point in space; (ii) the ‘tip’ of the instrument must track the reference indicated by the surgeon.
@@ III-B Virtual model control for laparoscopic surgery
Figure 1 provides a graphical representation of a virtual mechanism for laparoscopy. After gravity compensation, a ‘virtual instrument’ is virtually connected to the actual laparoscopic instrument attached to the robot, constraining and driving its motion. It is composed of the virtual instrument itself—a rigid body structure with a spherical joint located at the RCM and a co-located prismatic joint—and a set interface spring-damper pairs. By construction, axis of the virtual instrument always lies on the RCM. The interface springs pull the instrument towards the virtual instrument. At equilibrium, the instrument and the virtual instrument will be co-located,satisfying the RCM constraint. At the same time, the spring and damper between the end effector at the tip of the instrument and the reference position $\bm{r}$ will cause the end effector to track the surgeon’s motion.
Compared to PD based approaches, our approach does not require inverse kinematics, and therefore gracefully handles redundancy in the robot structure, producing an unconstrained ‘null space’ (see Figure 8, left). The controller is essentially identical for different robot morphologies. The spring placement allows us to stiffen the robot’s behaviour at at the RCM independently of the end-effector: compared to energy shaping, the mechanical analogy of the virtual mechanism makes design alterations intuitive, guided by our understanding of the physics of the system. The example on the right of Figure 8 shows the effect of replacing the linear springs between the robot and virtual instrument with saturating springs (described in Figure 4), when faced with a step disturbance force at the port. The saturating spring allows a compliant behaviour at the RCM for large disturbances, while maintaining the constraint for smaller forces. This could be used as an alternative approach to safety at the trocar: failing to a compliant mode rather than a stiff position hold mode, depending on the risk analysis. This design change is enabled by the virtual mechanism analogy.
To implement virtual model controllers no additional sensing beyond torque actuated joints with position/velocity feedback is required, and dynamic tuning/filtering can be achieved by changing inertance/damping of virtual instrument. Importantly, because the virtual mechanism is passive, the system is stable for many types of model uncertainty, and external interactions. If the actuators are capable of faithfully producing the desired control forces, then the system is stable regardless of the parameterization of the controller (although in practise there are upper limits on stiffness and damping that can be realized). This helps bridge the sim-to-real gap, as all models are uncertain to a degree, and additionally the system has many unmodified external interactions.
@@
The virtual mechanism presented here is based on the ‘Virtual Instrument’ of [38], but has a simplified structure. This is due to the tuning method discussed in the next section, which allows for more general virtual mechanisms than [38].
Refer to caption Refer to caption
Figure 8: In each of these figures, red indicates distance from the instrument axis to the RCM, and blue indicates distance from the instrument tip to the reference. [Top-Left] Null space motion: starting pose (solid) and middle/final pose (translucent) for a constant reference/rcm with a 1Nm torque applied at J1 for 12 seconds. [Bottom-Left] End-effector error remains low while moving in the null-space. [Right] Position error at the port and end-effector when a disturbance force is applied at the port. The disturbance force is either 5N (dashed lines) or 10N (solid lines). The top version has linear springs between the virtual-instrument and robot, the bottom has springs that saturate at 5N. We see similar behaviour at the port for smaller forces, but compliant behaviour for a large disturbance. In all cases the end-effector error remains low.
IV Optimal virtual mechanisms:
@@
IV-A Input/output performance
@@
In what follow we consider the robot as an open system, as shown in Figure 9. The robot is represented by the upper block, $R$, and the virtual model controller is represented by the lower block, $C$. Performance inputs ${\bm{w}}$ model exogenous perturbations to the robot that arise from the interaction with the working environment, such as contact forces. Performance outputs ${\bm{y}}$ capture the behavior of interest, that is, displacements and velocities (dynamically weighted) that are relevant to the task. The virtual parameters of the controller are represented by the vector ${\bm{\theta}}$, whose elements are greater than or equal to zero, modeling stiffness and damping coefficients, and inertia and geometric parameters such as length and weight of virtual links.
@@
Figure 9 represents a classical setting of linear robust control theory [48], whose objective is to design the controller to guarantee specific closed-loop performance from ${\bm{w}}$ to ${\bm{y}}$. In this paper, the controller is restricted to a family of (nonlinear) virtual mechanisms, parameterized by ${\bm{\theta}}$, and the goal is to find the parameters ${\bm{\theta}}$ that optimize a combined $\mathcal{L}_{2}$ and $\mathcal{L}_{\infty}$ performance from ${\bm{w}}$ to ${\bm{y}}$. Namely, considering the splitting of the signals ${\bm{w}}=({\bm{w}}_{1},{\bm{w}}_{2})$ and ${\bm{y}}=({\bm{y}}_{1},{\bm{y}}_{2})$, we want to optimize the aggregated cost
@@
[…]
@@
By optimizing over ${\bm{\theta}}$ in (13), we constrain the control action within a predefined set of virtual mechanisms while shaping the response of the controlled robot from ${\bm{w}}$ to ${\bm{y}}$, i.e. its closed-loop impedance ${\bm{w}}$ to ${\bm{y}}$. In this sense, (13) extends the classical framework of interconnection and damping assignment [3] to impedance control [16] and performance optimization.
IV-B Tractable optimization
The min-max form of (13) is a notoriously difficult type of optimization problem [49]. In contrast to linear control theory [48], (13) is far from tractable in the nonlinear and structured setting of virtual mechanisms, as it requires infinitely many probing signals ${\bm{w}}$. To solve this optimization, we need to build a sequence of approximations that leads to a well-defined, tractable optimization problem. We propose two approaches.
@@ IV-B Tractable optimization
$\mathcal{W}_{k\!+\!1}=\mathcal{W}_{k}\cup\{{\bm{w}}(\overline{\omega})\}\mbox{for }\overline{\omega}=\mathrm{argmax}_{\omega}\,L({\bm{\theta}}_{k},{\bm{w}}(\omega))\vspace{2mm}$ (15b)
@@
At the cost of suboptimality, these two approaches simplify the optimization since the inner maximization is now reduced to a finite set of signals (possibly growing). The outer minimization can be solved via gradient descent, as discussed in the next section. Even if suboptimal, Figures 11-13 show that a sufficiently dense set $\mathcal{W}$ for the sampling approach, or a few iteration steps for the adversarial one, provides neat performance results.
IV-C Solution via algorithmic differentiation
The minimization problem above can be solved by algorithmic differentiation, [24, 50]. The optimal vector ${\bm{\theta}}$ can be computed by gradient descent through ODE simulations. We make use of the Julia package SciMLSensitivity.jl [23], which integrates with a modular algorithmic differentiation ecosystem in Julia to provide differentiation rules for many ODE solvers in forward and backward modes. In this paper, we use automatic differentiation in the forward mode, which becomes slower to compute the gradient as the number of parameters in ${\bm{\theta}}$ scales, but is fast for a small number of parameters and can be used with minimal code adaptation.
@@ IV-C Solution via algorithmic differentiation
The gradient of ${\bm{x}}$ at a generic time $t$ is thus obtained by initializing (16) with a suitable selection of $\frac{\partial{\bm{x}}}{\partial{\bm{\theta}}}$ at time 00. Thus, to use automatic differentiation, we need to reformulate our optimization problem as a (differentiable) ODE of the form $\dot{{\bm{x}}}=f({\bm{x}},{\bm{\theta}})$, possibly paired to some readout function $h({\bm{x}}(t))$, where $h$ (differentiable) nonlinear mapping of the vector ${\bm{x}}$ at time $t$. Together, the ODE and the readout function must capture the complete dynamics of the robot and the virtual mechanism. They must also contain additional state variables to compute the aggregate cost $L$ in (13).
@@
[…]
@@
V Optimal virtual mechanisms:
@@
V-A Optimal PD control of a cart
We illustrate the controller optimization on the simplest example of a PD control of a cart. From Figure 10, the equations of cart and controller read

### v3

@@ III-C Virtual structures and implementation
Figure 5: Virtual model controller block diagram. $({\bm{F}}_{e},{\bm{z}}_{e})$ is an external port. $r$ represent external inputs to the controller. This is a general structure for any torque controller robot providing joint position and velocity feedback. In the case of the virtual instrument controller of Fig. 1, described in detail in Section IV-B, the ‘virtual structure’ comprises of the rotating joints, and the mass/inertia of the virtual instrument itself, while the springs and dampers between the virtual instrument and robot are ‘interface components’.
@@
If the many Jacobians need to be constructed to apply forces at a large number of points, these computations can form a significant fraction of the computational cost. To improve performance in future implementations, explicit computation of Jacobians can be avoided entirely for serial robots or tree-structured robots by recursive algorithms, which propagate kinematics/velocities forwards through the structure from the base, and forces backward from the tip(s), while also computing the required joint actuation $\bm{u}$}. For an example, refer to the the backwards pass of the Recursive Newton Euler Algorithm in [Featherstone2008].
@@
Although we limit our discussion to fully actuated robots, virtual model control is still possible for underactuated robots. The mechanism must in that case be designed such that the un-actuated joints are not required to emulate the virtual forces on the robot. Indeed, in Fig. 2, the force ${\bm{F}}_{1}$ does not require any actuation of the third joint. Likewise, in a quadruped, the underactuation is due to the lack of rigid connection to the ground. However, one could design a virtual mechanism acting between any part of the body/legs of the robot [Pratt1995b].
IV The design of virtual mechanisms
IV-A Virtual model control for reaching
The advantages of designing a controller as a virtual mechanisms can be illustrated by revisiting the classical control of a ‘reach’ motion, the basic primitive of any manipulator. Once gravity is compensated, the simplest virtual mechanism is given by a virtual spring and a virtual damper connected between the end effector and the target. This corresponds to the classical PD control in the task space [Lawrence1988]. It is straightforward to see that the adoption of nonlinear elements in this context enables important features. Consider a spring and a damper with nonlinear force-displacement characteristics
@@ IV-A Virtual model control for reaching
The more advanced exemplary virtual mechanism of Fig. 6 allows for refined control over smoothness and peak velocity of the reaching motion. Linear spring and dampers are now attached between the end effector and virtual moving cart. The cart is pulled by a constant force, experiences some drag due to friction, and has mass, which leads to a dynamic controller.
This design has several useful features. It separates the tasks of constraining the robot and driving motion along the line: the spring-damper components between the damper and the cart regulate the constraint, while forces/components applied to the cart regulate the reaching motion. By driving the cart with a constant force pointed in the direction of the final desired position, rather than as by a reference position that moves as a function of time, we achieve a safer form of path following. Consider an obstruction such as a person in the path of the robot: the end-effector and person collide and the cart continues to move forwards only until equilibrium is reached between the driving force and the constraint-spring. Thus, even for a stiff spring ensuring a strong constraint to the path, we can achieve compliant behaviour along the direction of motion. In addition the velocity of the robot end-effector can be regulated by tuning the friction applied to the cart: neglecting the robot’s friction, the steady state velocity can be determined by balancing the cart’s friction with the applied driving force. Finally, robot fast accelerations can be moderated by increasing the inertance of the cart. The cart also has the effect of filtering the driving force: a step change in driving force due to a sudden change of reference (or noise) results in smooth driving torques. For a numerical example of these effects, see Fig. 6.
@@
The above considerations illustrate how the design of virtual mechanisms carries significant insights on the control action. While equivalent behaviour can be achieved with impedance/force control, virtual mechanisms provide a framework for reasoning about controller design decisions, particularly in cases more complicated than this pedagogical example. By relying on physical intuition, we can make predictions about the robot reactions to obstacles, the velocity peak of the end effector, and the filtering effects of the controller dynamics. It is also easier to understand how the parameters of the virtual mechanism will affect the control performance. For example, it is straightforward to imagine that a larger cart mass will induce a stronger lowpass effect on the pulling force, as shown in Fig. 6. Similarly, a larger spring stiffness may lead to higher peak forces.
IV-B Virtual model control for laparoscopic surgery
Virtual model control appears particularly effective on complex tasks. Here we consider robotic surgery, specifically laparoscopic surgery, as shown in Fig. 7. The controller has two main objectives: (i) the instrument must pass through a small aperture in the patient’s body, often achieved by maintaining the tool over the so-called Remote Center of Motion (RCM), a fixed point in space; (ii) the ‘tip’ of the instrument must track the reference indicated by the surgeon.
@@ IV-B Virtual model control for laparoscopic surgery
Fig. 1 provides a graphical representation of a virtual mechanism for laparoscopy. After gravity compensation, a ‘virtual instrument’ is virtually connected to the actual laparoscopic instrument attached to the robot, constraining and driving its motion. It is composed of the virtual instrument itself—a rigid body structure with a spherical joint located at the RCM and a co-located prismatic joint—and a set interface spring-damper pairs. By construction, axis of the virtual instrument always lies on the RCM. The interface springs pull the instrument towards the virtual instrument. At equilibrium, the instrument and the virtual instrument will be co-located, satisfying the RCM constraint. At the same time, the spring and damper between the end effector at the tip of the instrument and the reference position $\bm{r}$ will cause the end effector to track the surgeon’s motion.
Compared to PD based approaches, our approach does not require inverse kinematics, and therefore gracefully handles redundancy in the robot structure, producing an unconstrained ‘null space’ (see Fig. 8). The controller is essentially identical for different robot morphologies. The spring placement allows us to stiffen the robot’s behaviour at the RCM independently of the end-effector: compared to energy shaping, the mechanical analogy of the virtual mechanism makes design alterations intuitive, guided by our understanding of the physics of the system. The plots of Fig. 9 shows the effect of replacing the linear springs between the robot and virtual instrument with saturating springs (described in Fig. 4), when faced with a step disturbance force at the RCM. The saturating spring allows a compliant behaviour at the RCM for large disturbances.
So long as the saturation force of the springs is larger than the standard required operational forces, typical behaviour is minimally affected. These springs could allow surgical staff to manually override the controller by pulling with a large enough force, and we use a similar design for robot assisted drilling in [Larby2025]. This design change is enabled by the virtual mechanism analogy.
To implement virtual model controllers no additional sensing beyond torque actuated joints with position/velocity feedback is required, and dynamic tuning/filtering can be achieved by changing inertance/damping of virtual instrument. Importantly, because the virtual mechanism is passive, the system is stable for many types of model uncertainty (parametric or dynamic perturbations preserves the passivity of the robot, and external interactions as long as the environment remains passive). If the actuators are capable of faithfully producing the desired control forces, then the system is stable regardless of the parameterization of the controller (although in practise there are upper limits on stiffness and damping that can be realized). This helps bridge the sim-to-real gap, as all models are uncertain to a degree, and additionally the system has many unmodified external interactions.
@@
The virtual mechanism presented here is based on the ‘Virtual Instrument’ presented in [Larby2022b], but has a simpler structure. This is due to the tuning method discussed in the next section, which allows for more general virtual mechanisms than [Larby2022b].
Figure 8: Null space motion from a starting pose (solid) and middle/final poses (translucent) for a constant end-effector/RCM position with a 1Nm torque applied at J1 for 12 seconds. End-effector and RCM error remain low ($1\text{\,}\mathrm{m}\mathrm{m}$\$) while moving in the null-space.
Figure 9: Tracking error when a disturbance force is applied at the RCM. The disturbance force is either 5N or 10N. The two graphs show results with linear springs between the virtual-instrument, or saturating springs that saturate at 10N. For small disturbance forces, behaviour is similar, but for large disturbances we see compliant behaviour for a large disturbance. In all cases the end-effector error remains low.
V Optimal virtual mechanisms:
@@
V-A Input/output performance
@@
In what follow we consider the robot as an open system, as shown in Fig. 10. The robot is represented by the upper block, $R$, and the virtual model controller is represented by the lower block, $C$. Performance inputs $\bm{w}$} model exogenous perturbations to the robot that arise from the interaction with the working environment, such as contact forces. Performance outputs $\bm{y}$} capture the behavior of interest, that is, displacements and velocities (dynamically weighted) that are relevant to the task. The virtual parameters of the controller are represented by the vector $\bm{\theta}$}, whose elements are greater than or equal to zero, modeling stiffness and damping coefficients, and inertia and geometric parameters such as length and weight of virtual links.
@@
Fig. 10 represents a classical setting of linear robust control theory [Zhou1996], whose objective is to design the controller to guarantee specific closed-loop performance from $\bm{w}$} to $\bm{y}$}. In this paper, the controller is restricted to a family of (nonlinear) virtual mechanisms, parameterized by $\bm{\theta}$}, and the goal is to find the parameters $\bm{\theta}$} that optimize a combined $\mathcal{L}_{2}$ and $\mathcal{L}_{\infty}$ performance from $\bm{w}$} to $\bm{y}$}. Namely, considering the splitting of the signals ${\bm{w}}=({\bm{w}}_{1},{\bm{w}}_{2})$ and ${\bm{y}}=({\bm{y}}_{1},{\bm{y}}_{2})$, we want to optimize the aggregated cost
@@
with the standard meaning $\parallel\!{\bm{w}}_{1}\!\!\parallel_{2}=\left(\int_{0}^{\infty}\!{\bm{w}}_{1}(t)^{T}{\bm{w}}_{1}(t)dt\right)^{\frac{1}{2}}$ and $\parallel\!{\bm{w}}_{2}\!\!\parallel_{\infty}=\sup_{t\geq 0}|{\bm{w}}_{2}(t)|$, assuming boundedness of the input signals in their respective domains.
@@
By optimizing over $\bm{\theta}$} in (19), we constrain the control action within a predefined set of virtual mechanisms while shaping the response of the controlled robot from $\bm{w}$} to $\bm{y}$}, i.e. its closed-loop impedance $\bm{w}$} to $\bm{y}$}. In this sense, (19) extends the classical framework of interconnection and damping assignment [Ortega2008] to impedance control [Hogan1984] and performance optimization.
V-B Tractable optimization
The min-max form of (19) is a notoriously difficult type of optimization problem [Hsieh2021]. In contrast to linear control theory [Zhou1996], (19) is far from tractable in the nonlinear and structured setting of virtual mechanisms, as it requires infinitely many probing signals $\bm{w}$}. To solve this optimization, we need to build a sequence of approximations that leads to a well-defined, tractable optimization problem. We propose two approaches.
@@ V-B Tractable optimization
$\mathcal{W}_{k\!+\!1}=\mathcal{W}_{k}\cup\{{\bm{w}}(\overline{\omega})\}\mbox{ for }\overline{\omega}=\mathrm{argmax}_{\omega}\,L({\bm{\theta}}_{k},{\bm{w}}(\omega))\vskip 5.69054pt$ (21b)
@@
At the cost of suboptimality, these two approaches simplify the optimization since the inner maximization is now reduced to a finite set of signals (possibly growing). The outer minimization can be solved via gradient descent, as discussed in the next section. Even if suboptimal, Figs. 12-14 show that a sufficiently dense set $\mathcal{W}$ for the sampling approach, or a few iteration steps for the adversarial one, provides neat performance results.
V-C Solution via algorithmic differentiation
The minimization problem above can be solved by algorithmic differentiation, [Rackauckas2021a, Ma2021]. The optimal vector $\bm{\theta}$} can be computed by gradient descent through ODE simulations. We make use of the Julia package SciMLSensitivity.jl [Rackauckas2021], which integrates with a modular algorithmic differentiation ecosystem in Julia to provide differentiation rules for many ODE solvers in forward and backward modes. In this paper, we use automatic differentiation in the forward mode, which becomes slower to compute the gradient as the number of parameters in $\bm{\theta}$} scales, but is fast for a small number of parameters and can be used with minimal code adaptation.
@@ V-C Solution via algorithmic differentiation
The gradient of $\bm{x}$} at a generic time $t$ is thus obtained by initializing (22) with a suitable selection of $\frac{\partial{\bm{x}}}{\partial{\bm{\theta}}}$ at time 0. Thus, to use automatic differentiation, we need to reformulate our optimization problem as a (differentiable) ODE of the form $\dot{{\bm{x}}}=f({\bm{x}},{\bm{\theta}})$, possibly paired to some readout function $h({\bm{x}}(t))$, where $h$ (differentiable) nonlinear mapping of the vector $\bm{x}$} at time $t$. Together, the ODE and the readout function must capture the complete dynamics of the robot and the virtual mechanism. They must also contain additional state variables to compute the aggregate cost $L$ in (19).
@@
  • For the $\mathcal{L}_{2}$ part, we replace the indefinite integral $\int_{0}^{\infty}\dots dt$, with the definite one $\int_{0}^{T}\dots dt$, where $0,T]\subset\mathbb{R}$ is the chosen simulation interval. Then, we compute this integral by introducing two new state variables, ${c}_{{\bm{w}}_{1}}$ and ${c}_{{\bm{y}}_{1}}$, and the ODE formulation
@@
VI Optimal virtual mechanisms:
@@
VI-A Optimal PD control of a cart
We illustrate the controller optimization on the simplest example of a PD control of a cart. From Fig. 11, the equations of cart and controller read